OpenStudy (anonymous):

7,19,37,61,91....... is this a sequence if yes mention the type of the sequence.

hartnn (hartnn):

Doesn't look like a sequence to me...

OpenStudy (anonymous):

my i call it hexagonal centred numbers sequence

hartnn (hartnn):

i just look that up. yes, thats correct (even though the '1' in the beginning is missing.)

hartnn (hartnn):

3n (n-1)+1

OpenStudy (agent0smith):

+12, +18, +24, +30 are the differences... which yes looks hexagonal: http://www.drking.org.uk/hexagons/misc/numbers.html

hartnn (hartnn):

yes, 2nd differences are constant +6 so the n'th term had to be quadratic.

OpenStudy (anonymous):

without 1 may it be hexagonal centred numbers ?

hartnn (hartnn):

yes, yes, just don't forget to mention that 'n' cannot take the value of 1 an = 3n(n-1)+1 for n=2,3,4....

OpenStudy (anonymous):

find 4th term of the give sequence an show me it is correct

hartnn (hartnn):

for 2st term, put n=2, so for 4th term put n=5 but thats not the usual way, so we need to edit the formula an = 3(n+1)n +1 now put n=4 for 4th term

OpenStudy (agent0smith):

Replace n with n+1 an = 3(n+1)(n)+1 for n=1,2,3...

hartnn (hartnn):

for 1st term**

OpenStudy (anonymous):

so what is the 4th term according to your formula

OpenStudy (agent0smith):

Replacing n with n+1 skips the first term, and yields 7,19,37,61,91

hartnn (hartnn):

4th term ----> n=4 a4 = 3*4*5+1 = 61

OpenStudy (anonymous):

replacing n with n+1 is right or wrong

hartnn (hartnn):

why that doubt ? we initially had 3n(n-1)+1 but n started from 2. to make it start from 1, we replaced 'n' by 'n+1' (yes, we can) so we get an = 3n(n+1)+1 , n=1,2,3....

OpenStudy (agent0smith):

Why do you keep doubting us? Why not check our formulas for yourself to see if they're valid...? if n=1, then n+1 = 2. Thus, replacing n with n+1 skips the first term (which would be 1).

OpenStudy (agent0smith):

^ since as hartnn said, "for 1st term, put n=2, so for 4th term put n=5"